OpenStudy (anonymous):

y varies directly as the square of x. When x = 1, y = 3. What is the value of y when x = 3?

OpenStudy (ghazi):

Y=X^2, now just substitute X=3 you'll get Y

OpenStudy (anonymous):

We need the equation y = kx in the general form. but we have y = kx^2 and substituting in the x = 4 and y = 32 value we get 32 = k((4)^2 32 = 16k 2 = k so the equation becomes y = 2x^2 if we want to find x when y = 18 we get 18 = 2x^2 9 = x^2 x = 3 or x = -3

OpenStudy (anonymous):

my options are: 1 9 27 81

OpenStudy (ghazi):

9

OpenStudy (anonymous):

could you explain why please?

OpenStudy (anonymous):

u can see in my ans

OpenStudy (ghazi):

just put X=3 and then take square of it...since Y=X^2 therefore it is just square of three

OpenStudy (anonymous):

taken from http://answers.yahoo.com/question/index?qid=20100924160042AAdzBNT

OpenStudy (anonymous):

ohh thanksss everyone!

OpenStudy (anonymous):

welcom